OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8>

OpenStudy (anonymous):

use this formula \[\cos(\theta)=\frac{u.v}{|u||v|}\]

OpenStudy (anonymous):

so cos =2(3)/-4(-8)?

OpenStudy (anonymous):

no take the dot product

OpenStudy (anonymous):

what's that?

OpenStudy (anonymous):

@NotSObright

OpenStudy (anonymous):

use your textbook learn it first its important or you cannot calculate angle

OpenStudy (anonymous):

i dont understand it in the textbooks i just need the steps to be explained easier @NotSObright